# It Cannot Be Right If It Was Written by AI: On Lawyers’ Preferences of Documents Perceived as Authored by an LLM vs a Human

### v1

@@ Abstract
Large Language Models (LLMs) enable a future in which certain types of legal documents may be generated automatically. This has a great potential to streamline legal processes, lower the cost of legal services, and dramatically increase access to justice. While many researchers focus their efforts on proposing and evaluating LLM-based applications supporting tasks in the legal domain, there is a notable lack of investigations into how legal professionals perceive content if they believe it has been generated by an LLM. Yet, this is a critical point as over-reliance or unfounded skepticism may influence whether such documents bring about appropriate legal consequences. This study is the necessary analysis in the context of the ongoing transition towards mature generative AI systems. Specifically, we examined whether the perception of legal documents’ by lawyers (n=75) varies based on their assumed origin (human-crafted vs AI-generated). The participants evaluated the documents focusing on their correctness and language quality. Our analysis revealed a clear preference for documents perceived as crafted by a human over those believed to be generated by AI. At the same time, most of the participants are expecting the future in which documents will be generated automatically. These findings could be leveraged by legal practitioners, policy makers and legislators to implement and adopt legal document generation technology responsibly, and to fuel the necessary discussions into how legal processes should be updated to reflect the recent technological developments.
@@ 1 Introduction
This paper analyzes the differences in lawyers’ perceptions of documents believed to be either generated by AI or authored by humans. The aim of this study is to react to the revolution brought about by the recent advances in Large Language Models (LLMs), and provide crucial insights necessary for responsible design, development, and adoption of LLM-powered applications focused on automated generation of legal documents. Should it be known it is LLM-generated, potential distortions in the perception of a legal document present a serious ethical concern. For example, it would be highly undesirable if a judge denies a rightful claim because they perceive the arguments in a complaint as less persuasive, knowing the document has been automatically generated. To investigate this pressing concern, we had a large group of lawyers (n=75) evaluate documents marked as AI-generated or human-crafted in terms of their correctness and language quality. Further, we surveyed the participants on their beliefs on the ability of LLMs generating legal documents in the future.
@@
ChatGPT [^1] OpenAI: ChatGPT. Available at: https://chat.openai.com/ [Accessed 2024-03-16] was launched on November 30, 2022, and it immediately sensitised the general public to the capabilities of LLMs to write fluent texts and lead conversations in natural language. While the foundational GPT-3 model has been available since 2020 [12], it was the accessibility of the ChatGPT service that made legal practitioners, educators, and scholars alike engage in heated discussions as to how the legal profession may change in the near future. There have been numerous examples of applications of the technology to tasks traditionally reserved exclusively for legal experts. For example, Perlman [77] claims that ChatGPT requires re-imagining how to obtain legal services and prepare lawyers for their careers—he does so in a human-authored abstract to a scholarly article automatically generated with ChatGPT. Katz et al [49] tested GPT-4 on the Uniform Bar Examination (UBE), reporting the system passed the exam. Blair-Stanek et al [7] show how even smaller LLMs can achieve near perfect performance on basic tasks involving legal texts, if they are fine-tuned. Such use cases hint at future applications of LLMs in providing legal services and increasing access to justice, namely answering legal questions, providing legal information, and importantly drafting legal documents.
@@
The potential for automation, enhancement or support through LLMs is growing. Due to user-friendly access to LLMs through applications and interfaces (e.g., ChatGPT, Bard, Copilot), the experimentation with AI-based tools is available to masses desiring to become early adopters of the disruptive technology. However, the challenge of the wide use of LLMs lies in more than just its ability to achieve high (even human-like) accuracy when engaging in various tasks. If the output of LLMs is not acknowledged or perceived as accurate by the relevant actors, it may become inconsequential even if objectively accurate. The reception and perception of LLMs’ output, which is tied to various societal and psychological aspects of communication, is a pressing and understudied issue.
@@
Existing interdisciplinary research suggests generally negative sentiment towards technology, ranging from hesitancy [25] to aversion [44, 14]. Studies focusing on technology acceptance in the legal domain are scarce [99, 4, 74]. To our knowledge, no study directly focuses on examining lawyers’ preferences towards documents perceived as generated by AI (specifically by ChatGPT).
@@
To investigate if and how lawyers’ perception of a legal document changes depending on whether it is believed to be human-crafted as opposed to AI-generated, we analyzed the following research questions:
@@
    What are lawyers’ attitudes and beliefs regarding the potential of LLMs to generate documents automatically in the future?
Our work contributes to AI & Law research in the following ways. To our knowledge, this is the first comprehensive study that investigates differences in lawyers’ perceptions of documents believed to be crafted by a human as opposed to LLM-generated. It focuses on the correctness of the documents, i.e., their potential to emulate the desired legal outcomes, and their language properties, i.e., readability or absence of grammatical/stylistic errors.
The remainder of this article is structured as follows. Section 2 contains an overview of the related work focused on the perception of AI-generated documents, and on applications of LLMs in the legal domain. Section 3 presents the experimental design, focusing on the documents and the survey used in this study, and the participating subjects. In Section 4, we present the results of quantitative and qualitative analysis of the data collected from the participants. Section 5 engages in a discussion of our findings. In Section 6, we draw implications of our research for legal practice. Section 7 focuses on the limitation of the study. Finally, Section 8 summarizes the article and outlines the potential paths for future work.
@@ 2 Related Work
Experiments by Bubeck et al [13] and a survey by Naveed et al [67] have shown that LLMs offer possibilities and promises in various domains, including law. Law is based mainly on written language and is often lauded for its overwhelming production of information and documents. With the ever-increasing interest in NLP in the legal domain noted by Katz et al [48], significant advances in law-related NLP research were brought by pre-trained Transformed-based Language Models (TLMs) reviewed by Greco and Tagarelli [30]. The subsequent introduction of LLMs to the public fueled significant interest in its law-related applications noted by Lai et al [53].
@@
    Perception of AI-generated Content in Various Domains (Subsection 2.1) is focused on the crucial underlying issue of the perception of AI-generated content by public and other relevant actors. The issue must be separated from the performance demonstrated by LLMs when engaging in various tasks.
@@
    Automated Content Generation by LLMs in the Legal Domain (Subsection 2.2) focuses on related work reporting the use of LLMs to create summaries, translations and answers to legal questions.
@@
    The Use of LLMs in the Legal Domain (Subsection 2.3) summarizes related work reporting the use of LLMs for various tasks, such as legal reasoning, legal research, access to justice, annotation, and legal judgment prediction.
The related work clearly demonstrates a need for empirical studies regarding the perception of AI-generated documents in the legal domain. The overview below contains dozens of papers on supporting, enhancing or automating a plethora of tasks inherently perceived as requiring human attention. While the authors often report lower-than-human performance, GPT-4 significantly improves the already impressive capabilities. As a result, the use of LLMs is likely to penetrate every aspect of law. The remarkable breadth of the research conducted in the short time since November 2022 demonstrates the inevitability of AI, even in the legal domain. As such, the issue of the perception of AI-generated content in the legal domain becomes even more pressing.
@@ 2.1 Perception of AI-generated Content in Various Domains
Increased aversion is reported when algorithms are involved in moral decision-making, particularly in domains deemed morally significant such as medicine and law Bigman and Gray [5]. Laakasuo et al [52] demonstrated that decisions made by human-like robots are perceived as less moral compared to robots without resemblance to humans. Interestingly, AI-mediated communication provides flexibility in interpersonal interactions, as users are inclined to attribute part of the responsibility for negative feelings to the AI [36].
Lower evaluations of AI-generated content have been reported across various domains, including Airbnb profiles [40], emails [58], artworks [78], music [86], translations [2], news articles [93] or health prevention messages [57]. The level of aversion to algorithms differs between tasks perceived as subjective and objective, with perceived task objectivity leading to greater acceptance [15]. In healthcare, people tend to prefer human practitioners over AI-based technologies [65], expressing concerns over insufficient personalization of care and incompetence [60]. Wang et al [95] emphasized the necessity of incorporating essential soft skills and core principles, such as professionalism, explainability, and empathy, into GPT-based tools for medical consultations to overcome the aversion. Understanding the impact of AI-generated content on interactions within the legal domain is of utmost importance to ensure safe deployment of AI-based tools and applications.
While the algorithms may outperform humans in specific tasks, the post-error dynamic differs significantly. Dietvorst et al [23] found that people lost confidence more quickly in algorithms after witnessing the same error made by both algorithmic and human forecasters. Given the inevitability of mistakes and errors, the aspect is particularly significant especially in the legal domain, where mistakes and errors carry serious consequences for welfare or even freedom.
Surveys have highlighted issues with the factuality of generated messages [94]. LLMs are susceptible to hallucinations, which poses a significant challenge for further deployment in the legal [16] or financial domains [45]. Information asymmetry is crucial, as those who stand to gain the most are also exposed to the most significant risks associated with hallucinations [20]. Individuals with lower literacy and education levels are at a higher risk of consuming unreliable content [76]. Dahlkemper et al [21] reported that even university students with limited prior knowledge of specific issues related to their domain of expertise find evaluating the factuality of ChatGPT-generated answers challenging. These findings may manifest themselves in the legal domain as more AI-based tools are deployed to support various legal tasks without properly understanding the related risks.
@@ 2.2 Automated Content Generation by LLMs in Legal Domain
Summarization Fine-tuning summarization models for niche domains may be prohibitively costly, which makes the use of LLMs (especially ChatGPT) appealing, particularly in the health domain [85, 90], and will most probably extend to the law as well. LLMs perform on par with human written summaries in news summarization [101, 28]. Deroy et al [22] explored using pre-trained abstractive summarization models and LLMs for summarization in the legal domain. They acknowledged the limitations of fully automatic approaches and proposed a human-in-the-loop approach with legal experts monitoring the quality of outcomes. LLMs have been utilized to create court decision summaries, contributing to enhanced public trust in judicial outcomes [1]. AI-generated summaries offer increased accessibility and lower the cognitive effort needed for understanding, especially for non-experts. The logical progression includes summarising legal and regulatory documents akin to transformer-based summarization methods [50]. Ramprasad et al [79] have reported efforts in this direction, focusing on the factuality of zero-shot summaries across three domains, including legal bills. Additionally, Gesnouin et al [26] introduced LLaMandement, a fine-tuned model for generating neutral summaries of legislative proposals. The performance of LLM-based summarization models presents a significant potential benefit to the legal community. However, the perception of its intended recipients remains largely unaddressed.
@@
Translation ChatGPT demonstrated good translation quality when powered by GPT-3.5 [47] or GPT-4 [43]. Vieira et al [92] warned about the influence of machine translation in the legal domain, highlighting its under-researched impact on decision-making in critical legal situations. Studies on the use of LLMs for legal translation are limited. Mahapatra et al [61] developed a parallel corpus of legislation in Indian languages and evaluated various machine translation methods, including LLMs. They noted the mediocre performance of LLMs compared to other methods. Additionally, Briva-Iglesias et al [11] found that Google Translate outperforms LLMs, but GPT-4 is rated higher by evaluators for contextual adequacy and translation fluency. Reported performance hints at significant potential for future development and deployment. However, the translation of legal texts adds the language barrier on top of challenges associated with (lack of) legal knowledge. The understudied perception of accuracy may play a significant role.
@@
Legal Question Answering Given the awareness of ChatGPT among both the general and professional public, its use to generate answers to domain-specific questions is obvious. Before the introduction of ChatGPT, Metzler et al [64] proposed a combination of multi-task learning and zero- and few-shot learning to enhance question-answering capabilities. The performance of LLMs in domain-specific exams attracts significant research attention. In the health domain, Kung et al [51] reported ChatGPT achieving or surpassing the passing threshold in three exams without fine-tuning the general model. These findings show promise in using ChatGPT in medical education and potentially clinical decision-making. Bommarito et al [8] analyzed LLMs’ performance in an examination developed by the American Institute of Certified Public Accountants containing questions related to law, finance, accounting, technology, and ethics. The research showed that GPT-3.5 is approaching human-level performance in Remembering & Understanding, and Application skill portions of the exam. These results led the authors to conclude that LLMs have transformative potential in the future of knowledge work. However, Shen et al [87] observed varying levels of ability of ChatGPT across different domains with significant underperformance in legal and science-related questions. Choi et al [18] investigated the use of GPT-3.5 in answering final exams in four law school courses at the University of Minnesota, noting promising but uneven performance. Overall, its question-answering performance was deemed mediocre but sufficient to earn a J.D. degree eventually. Katz et al [49] reported significant improvement with GPT-4, outperforming prior models and even humans in five out of seven categories of questions. Similarly Nay et al [69] noted improved accuracy of GPT-4 over previous models in answering multiple-choice questions about US tax law. On the other hand, Martínez [63] evaluated GPT-4’s UBE performance, and criticized the reported values as potentially harmful in misrepresenting the actual capability of LLMs to truthfully and accurately answer legal questions. Tan et al [89] compared answers generated by JusticeBot, an expert tool focused on landlord-tenant disputes, and ChatGPT. They noted ChatGPT’s lack of precision in providing legal information directly to laypeople. However, they suggested that combining its capabilities with expert systems may enhance question-answering performance and improve the accessibility of legal information.

### v2

@@ Abstract
Large Language Models (LLMs) enable a future in which certain types of legal documents may be generated automatically. This has a great potential to streamline legal processes, lower the cost of legal services, and dramatically increase access to justice. While many researchers focus on proposing and evaluating LLM-based applications supporting tasks in the legal domain, there is a notable lack of investigations into how legal professionals perceive content if they believe an LLM has generated it. Yet, this is a critical point as over-reliance or unfounded scepticism may influence whether such documents bring about appropriate legal consequences. This study is the necessary analysis of the ongoing transition towards mature generative AI systems. Specifically, we examined whether the perception of legal documents’ by lawyers and law students (n=75) varies based on their assumed origin (human-crafted vs AI-generated). The participants evaluated the documents, focusing on their correctness and language quality. Our analysis revealed a clear preference for documents perceived as crafted by a human over those believed to be generated by AI. At the same time, most participants expect the future in which documents will be generated automatically. These findings could be leveraged by legal practitioners, policymakers, and legislators to implement and adopt legal document generation technology responsibly and to fuel the necessary discussions on how legal processes should be updated to reflect recent technological developments.
@@ 1 Introduction
This paper analyzes the differences in lawyers’ and law students’ perceptions of documents believed to be either generated by AI or authored by humans. The aim of this study is to react to the revolution brought about by the recent advances in Large Language Models (LLMs), and provide crucial insights necessary for responsible design, development, and adoption of LLM-powered applications focused on automated generation of legal documents. Should it be known that documents are LLM-generated, potential distortions in perception present a serious concern. For example, it would be highly undesirable if a judge denies a rightful claim because they perceive the arguments in a complaint as less persuasive, knowing the document has been automatically generated. To investigate this pressing concern, we had a large group of lawyers and law students (n=75) evaluate documents labeled as AI-generated or human-crafted in terms of correctness and language quality. Further, we surveyed the participants on their beliefs on the ability of LLMs to generate legal documents automatically in the future.
@@
ChatGPT [^1] OpenAI: ChatGPT. Available at: https://chat.openai.com/ [Accessed 2024-08-23] was launched on November 30, 2022, and it immediately sensitized the general public to the capabilities of LLMs to write fluent texts and lead conversations in natural language. While the foundational GPT-3 model has been available since 2020 [12], it was the accessibility of the ChatGPT service that made legal practitioners, educators, and scholars alike engage in heated discussions as to how the legal profession may change in the near future. There have been numerous examples of applications of the technology to tasks traditionally reserved exclusively for legal experts. For example, Perlman [80] claims that ChatGPT requires re-imagining how to obtain legal services and prepare lawyers for their careers—he does so in a human-authored abstract to a scholarly article automatically generated with ChatGPT. Katz et al [51] tested GPT-4 on the Uniform Bar Examination (UBE), reporting the system passed the exam. Blair-Stanek et al [7] show how even smaller LLMs can achieve near-perfect performance on basic tasks involving legal texts if they are fine-tuned. Such use cases hint at future applications of LLMs in providing legal services and increasing access to justice, namely answering legal questions, providing legal information, and, importantly, drafting legal documents.
@@
The potential for automation, enhancement or support through LLMs is growing. Due to user-friendly access to LLMs through applications and interfaces (e.g., ChatGPT, Bard, Copilot), the experimentation with AI-based tools is available to masses desiring to become early adopters of the disruptive technology. However, the challenge of the wide use of LLMs lies in more than just their ability to achieve high (even human-like) accuracy when engaging in various tasks. If the output of LLMs is not acknowledged or perceived as accurate by the relevant actors, it may become inconsequential even if objectively accurate. The reception and perception of LLMs’ output, which is tied to various societal and psychological aspects of communication, is a pressing and understudied issue.
@@
Existing interdisciplinary research suggests generally negative sentiment towards technology, ranging from hesitancy [27] to aversion [46, 14]. Studies focusing on technology acceptance in the legal domain are scarce [103, 4, 77]. To our knowledge, no study directly focuses on examining lawyers’ and law students’ preferences towards documents perceived as generated by AI (specifically by ChatGPT).
@@
To investigate if and how lawyers’ and law students’ perceptions of a legal document change depending on whether it is believed to be human-crafted as opposed to AI-generated, we analyzed the following research questions:
@@
    What are their attitudes and beliefs regarding the potential of LLMs to generate documents automatically in the future?
Our work contributes to AI & Law research in the following ways. To our knowledge, this is the first comprehensive study investigating differences in lawyers’ perceptions of documents believed to be crafted by a human as opposed to LLM-generated. It focuses on the correctness of the documents, i.e., their potential to emulate the desired legal outcomes, and their language properties, i.e., readability or absence of grammatical/stylistic errors.
The remainder of this article is structured as follows. Section 2 contains an overview of the related work focused on the perception of AI-generated documents and applications of LLMs in the legal domain. Section 3 presents the experimental design, focusing on the documents, the survey used in this study, and the participating subjects. In Section 4, we present the quantitative and qualitative analysis results of the data collected from the participants. Section 5 discusses our findings. In Section 6, we draw implications of our research for legal practice. Section 7 focuses on the limitations of the study. Finally, Section 8 summarizes the article and outlines the potential paths for future work.
To facilitate much-needed replications of our study, we release the documents and the survey used in our experiments in an accompanying online repository. [^2] See https://github.com/jakubharasta/AI_vs_Human_LegalDocs
@@ 2 Related Work
Experiments by Bubeck et al [13] and a survey by Naveed et al [70] have shown that LLMs offer possibilities and promises in various domains, including law. Law is based mainly on written language and is often lauded for its overwhelming production of information and documents. With the ever-increasing interest in NLP in the legal domain noted by Katz et al [50], significant advances in law-related NLP research were brought by pre-trained Transformer-based Language Models (TLMs) reviewed by Greco and Tagarelli [32]. The subsequent introduction of LLMs to the public fueled significant interest in its law-related applications, as noted by Lai et al [55].
@@
    Perception of AI-generated Content in Various Domains (Subsection 2.1) is focused on the crucial underlying issue of the perception of AI-generated content by the public and other relevant actors. The issue must be separated from the performance demonstrated by LLMs when engaging in various tasks.
@@
    Automated Content Generation by LLMs in Legal Domain (Subsection 2.2) focuses on related work reporting the use of LLMs to create summaries, translations and answers to legal questions.
@@
    The Use of LLMs in Legal Domain (Subsection 2.3) summarizes related work reporting the use of LLMs for various tasks, such as legal reasoning, legal research, access to justice, annotation, and legal judgment prediction.
The related work demonstrates a need for empirical studies focusing on the perception of AI-generated documents in the legal domain. The overview below contains dozens of papers on supporting, enhancing or automating a plethora of tasks inherently perceived as requiring human attention. While the authors often report lower-than-human performance, GPT-4 significantly improves the already impressive capabilities. As a result, the use of LLMs is likely to penetrate every aspect of law. The remarkable breadth of the research conducted in the short time since November 2022 demonstrates the inevitability of AI, even in the legal domain. As such, the under-studied issue of the perception of AI-generated content in the legal domain becomes even more pressing.
@@ 2.1 Perception of AI-generated Content in Various Domains
Increased aversion is reported when algorithms are involved in moral decision-making, particularly in domains deemed morally significant, such as medicine and law [5]. Laakasuo et al [54] demonstrated that decisions made by human-like robots are perceived as less moral compared to robots without resemblance to humans. Interestingly, AI-mediated communication provides flexibility in interpersonal interactions, as users are inclined to attribute part of the responsibility for negative feelings to the AI [38].
Lower evaluations of AI-generated content have been reported across various domains, including Airbnb profiles [42], emails [60], artworks [81], music [89], translations [2], news articles [97] or health prevention messages [59]. The level of aversion to algorithms differs between tasks perceived as subjective and objective, with perceived task objectivity leading to greater acceptance [15]. In healthcare, people tend to prefer human practitioners over AI-based technologies [68], expressing concerns over personalization of care and competence [62]. Wang et al [99] emphasized the necessity of incorporating essential soft skills and core principles, such as professionalism, explainability, and empathy, into GPT-based tools for medical consultations to overcome the aversion. Understanding the impact of AI-generated content on interactions within the legal domain is of utmost importance to ensure the safe deployment of AI-based tools and applications.
While the algorithms may outperform humans in specific tasks, the post-error dynamic differs significantly. Dietvorst et al [25] found that people lost confidence more quickly in algorithms after witnessing the same error made by both algorithmic and human forecasters. Given the inevitability of mistakes and errors, the aspect is particularly significant in the legal domain, where mistakes and errors have severe consequences for welfare or freedom.
Surveys have highlighted issues with the factuality of generated messages [98]. LLMs are susceptible to hallucinations, which poses a significant challenge for further deployment in the legal [16, 63] or financial domains [47]. Information asymmetry is crucial, as those who stand to gain the most are also exposed to the most significant risks associated with hallucinations [21]. Individuals with lower literacy and education levels are at a higher risk of consuming unreliable content [79]. Dahlkemper et al [22] reported that even university students with limited prior knowledge of specific issues related to their domain of expertise find evaluating the factuality of ChatGPT-generated answers challenging. These findings may manifest in the legal domain as more AI-based tools are deployed to support various legal tasks without adequately understanding the related risks.
@@ 2.2 Automated Content Generation by LLMs in Legal Domain
Summarization Fine-tuning summarization models for niche domains may be prohibitively costly, which makes the use of LLMs (especially ChatGPT) appealing, particularly in the health domain [88, 93], and will most probably extend to the law as well. LLMs perform on par with human written summaries in news summarization [105, 30]. Deroy et al [23, 24] explored using pre-trained abstractive summarization models and LLMs for summarization in the legal domain. They acknowledged the limitations of fully automatic approaches and proposed a human-in-the-loop approach with legal experts monitoring the quality of outcomes, addressing hallucinations and inconsistencies. LLMs have been utilized to create court decision summaries, contributing to enhanced public trust in judicial outcomes [1]. AI-generated summaries offer increased accessibility and lower the cognitive effort needed for understanding, especially for non-experts. The logical progression of using LLMs includes summarizing legal and regulatory documents akin to transformer-based summarization methods [52]. Ramprasad et al [82] have reported efforts in this direction, focusing on the factuality of zero-shot summaries across three domains, including legal bills. Additionally, Gesnouin et al [28] introduced LLaMandement, a fine-tuned model for generating neutral summaries of legislative proposals. The performance of LLM-based summarization models presents a significant potential benefit to the legal community. However, the perception by its intended recipients remains largely unaddressed.
@@
Translation ChatGPT demonstrated good translation quality when powered by GPT-3.5 [49] or GPT-4 [45]. Vieira et al [96] warned about the influence of machine translation in the legal domain, highlighting its under-researched impact on decision-making in critical legal situations. Studies on the use of LLMs for legal translation are limited. Mahapatra et al [64] developed a parallel corpus of legislation in Indian languages and evaluated various machine translation methods, including LLMs. They noted the mediocre performance of LLMs compared to other methods. Additionally, Briva-Iglesias et al [11] found that Google Translate outperforms LLMs, but GPT-4 is rated higher by evaluators for contextual adequacy and translation fluency. Reported performance hints at significant potential for future development and deployment. However, translating legal texts adds the language barrier to challenges associated with (lack of) specialized legal knowledge. The understudied perception of accuracy may play a significant role.
@@
Legal Question Answering Given the awareness of ChatGPT among both the general and professional public, its use to generate answers to domain-specific questions is obvious. Before the introduction of ChatGPT, Metzler et al [67] proposed a combination of multi-task learning and zero- and few-shot learning to enhance question-answering capabilities. The performance of LLMs in domain-specific exams attracts significant research attention. In the health domain, Kung et al [53] reported ChatGPT achieving or surpassing the passing threshold in three exams without fine-tuning the general model. These findings show promise in using ChatGPT in medical education and potentially clinical decision-making. Bommarito et al [8] analyzed LLMs’ performance in an examination developed by the American Institute of Certified Public Accountants containing questions related to law, finance, accounting, technology, and ethics. The research showed that GPT-3.5 is approaching human-level performance in Remembering & Understanding, and Application skill portions of the exam. These results led the authors to conclude that LLMs have transformative potential in the future of knowledge work. However, Shen et al [90] observed varying levels of ability of ChatGPT across different domains with significant underperformance in legal and science-related questions. Choi et al [18] investigated the use of GPT-3.5 in answering final exams in four law school courses at the University of Minnesota, noting promising but uneven performance. Overall, its question-answering performance was deemed mediocre but sufficient to earn a J.D. degree eventually. Katz et al [51] reported significant improvement with GPT-4, outperforming prior models and even humans in five out of seven categories of questions. Similarly, Nay et al [72] noted improved accuracy of GPT-4 over previous models in answering multiple-choice questions about U.S. tax law. On the other hand, Martínez [66] evaluated GPT-4’s UBE performance and criticized the reported values as potentially harmful in misrepresenting the actual capability of LLMs to truthfully and accurately answer legal questions. Tan et al [92] compared answers generated by JusticeBot, an expert tool focused on landlord-tenant disputes, and ChatGPT. They noted ChatGPT’s lack of precision in providing legal information directly to laypeople. However, they suggested that combining its capabilities with expert systems may enhance question-answering performance and improve the accessibility of legal information.